# Stability of $L^{p}$ Dirichlet solvability under small bi-Lipschitz transformations of domains

### v1

@@ Appendix B Solvability of the $L^{p}$ Dirichlet problem under $C^{1,\alpha}$ diffeomorphism of convex domains
(B.2) $\|\nabla u\|_{L^{\infty}(\Omega\cap B_{1/2})}\leq C\|u\|_{L^{\infty}(\Omega\cap B_{1})},$
@@
Clearly, $A_{\Phi}$ also satisfies the uniform ellipticity condition and $C^{\alpha_{1}}$ continuous (since $\Phi$ is a $C^{1,\alpha}$ diffeomorphism), with $\alpha_{1}=\min\{\alpha,\beta\}>0$. Hence, the desired estimate (B.2) is equivalent to
(B.3) $\|\nabla v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1/2}))}\leq C\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}.$
But now $\Phi^{-1}(\Omega\cap B_{1})=\{(x,t):t>\phi(x)\}\cap\Phi^{-1}(B_{1})$. By assumption, the part $\{(x,t):t=\phi(x)\}\cap\Phi^{-1}(B_{1})$ is convex. The gradient estimate (B.3) over convex boundaries is more or less well-known; see [GT83, Chapter 14.2] for elliptic operator in nondivergence form. We provide a proof for our elliptic operator in divergence form for the reader’s convenience.
@@
(B.4) $|v(0,t)|\leq Ct\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}$
This is proved by the comparison principle. We construct a barrier function in the following way. Let $M=\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}$. Let $w$ be a barrier function satisfying $L_{\Phi}(w)=0$ in $\{t>0\}\cap\Phi^{-1}(B_{1})$, $w=0$ on $\{t=0\}\cap\Phi^{-1}(B_{1})$ and $w=M$ on $\{t>0\}\cap\partial\Phi^{-1}(B_{1})$. Clearly, $0\leq w\leq M$ in $\{t>0\}\cap\Phi^{-1}(\Phi)$. The key fact we will use from the convexity of $\phi(x)$ is that $\{t>\phi(x)\}\cap\Phi^{-1}(B_{1})\subset\{t>0\}\cap\Phi^{-1}(B_{1})$ (due to the fact $\phi(x)\geq 0$). Therefore, we have $|v|=0\leq w$ on the convex part of the boundary $\{t=\phi(x)\}\cap\Phi^{-1}(B_{1})$ and $|v|\leq M=w$ on the rest of the boundary $\{t>\phi(x)\}\cap\partial\Phi^{-1}(B_{1})$. By the comparison principle, we have $|v|\leq w$ in $\{t>\phi(x)\}\cap\Phi^{-1}(B_{1})$. By the Schauder estimate of $w$ over a flat boundary $\{t=0\}$, we have $|w(x,t)|\leq CtM$ for any $(x,t)\in\{t>0\}\cap\Phi^{-1}(B_{3/4})$. This particularly implies (B.4).
Finally, by applying the previous argument for (B.4) to all the boundary points on $\{t=\phi(x)\}\cap\Phi^{-1}(B_{3/4})$, we actually have $|v(X)|\leq C\delta(X)M$, where $\delta(X)$ is the distance from $X$ to the boundary $\{t=\phi(x)\}$. Hence, by the interior Schauder estimate, for any $X\in\{t>\phi(x)\}\cap\Phi^{-1}(B_{1/2})$, we have
@@
This ends the proof of (B.3). ∎
Remark B.5.
@@
(B.7) $\left(\frac{1}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k(y)^{q}d\sigma(y)\right)^{1/q}\leq\frac{C}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k(y)d\sigma(y)=C\frac{\omega(\Delta(X,r))}{\sigma(X,r)}$
@@
by Lemma 2.11, where $G$ is the Green function for $\Delta$ with pole at infinity. Apply Lemma B.1 to $u=G$ in $B(X,r_{0})\cap\Omega$, we get that (for $s>0$ sufficiently small)
(B.8) $\frac{G(y^{\prime},s+\phi(y^{\prime}))}{s}\leq\frac{C\left\|G\right\|_{L^{\infty}(B(X,r)\cap\Omega)}}{r},$
@@
for all $r>r_{0}$. Combing the local estimate with the large-scale Lipschitz estimate, we obtain (B.8) for all $r>s>0$. By the Bourgain estimate (see e.g. [JK82a, Lemma 4.4]), one has $\left\|G\right\|_{L^{\infty}(B(X,2r)\cap\Omega)}\leq G(A_{X,r})$, where $A_{X,r}$ is a corkscrew point relative to $X$ and $r$, that is, $A_{X,r}\in B(X,Mr)\cap\Omega$ satisfies $\operatorname{dist}(A_{X,r},\partial\Omega)\approx|A_{X,r}-X|\approx r$. By Lemma 2.11 again, we have that $\frac{G(A_{X,r})}{r}\approx\frac{\omega(\Delta(X,r))}{\sigma(X,r)}$. Putting everything together, one has
@@
(B.9) $\left(\frac{1}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k^{X_{0}}(y)^{q}d\sigma(y)\right)^{1/q}\leq\frac{C}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k^{X_{0}}(y)d\sigma(y)=C\frac{\omega^{X_{0}}(\Delta(X,r))}{\sigma(X,r)}$
@@
where $g_{X}$ is convex with Lipschitz constant at most $M$. By a covering argument, it suffices to show (B.9) for $0<r<\min\left\{r_{0}/4,\operatorname{dist}(X_{0},\partial\Omega)/4\right\}$. Since for any $X\in\partial\Omega$ and $0<r<\min\left\{r_{0}/4,\operatorname{dist}(X_{0},\partial\Omega)/4\right\}$, $G(\cdot,X_{0})$ is a solution to $\Delta u=0$ in $B(X,2r)\cap\Omega$ with $G(\cdot,X_{0})=0$ on $\partial\Omega$, one can apply (B.1) to $G(\cdot,X_{0})$ in $B(X,r)\cap\Omega$ and the rest of the argument goes as in Case 1. ∎
Remark B.10.
In the case of an unbounded Lipschitz graph domain considered in Theorem B.6, if $\phi_{1}=0$, then $\Omega$ is convex. Then $\Phi_{x_{0}}$ can be taken as identity transform and $g_{x_{0}}(x)=\phi_{0}(x)$ for any $x_{0}\in\partial\Omega$.
Theorem B.11.
Under the same assumptions as Theorem (B.6), the $L^{p}$ Regularity problem $(R)_{p}$ for the Laplace operator is solvable in $\Omega$ for any $1<p<\infty$.

### v2

@@ Appendix C Solvability of the $L^{p}$ Dirichlet problem under $C^{1,\alpha}$ diffeomorphism of convex domains
(C.2) $\|\nabla u\|_{L^{\infty}(\Omega\cap B_{1/2})}\leq C\|u\|_{L^{\infty}(\Omega\cap B_{1})},$
@@
Clearly, $A_{\Phi}$ also satisfies the uniform ellipticity condition and $C^{\alpha_{1}}$ continuous (since $\Phi$ is a $C^{1,\alpha}$ diffeomorphism), with $\alpha_{1}=\min\{\alpha,\beta\}>0$. Hence, the desired estimate (C.2) is equivalent to
(C.3) $\|\nabla v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1/2}))}\leq C\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}.$
But now $\Phi^{-1}(\Omega\cap B_{1})=\{(x,t):t>\phi(x)\}\cap\Phi^{-1}(B_{1})$. By assumption, the part $\{(x,t):t=\phi(x)\}\cap\Phi^{-1}(B_{1})$ is convex. The gradient estimate (C.3) over convex boundaries is more or less well-known; see [GT83, Chapter 14.2] for elliptic operator in nondivergence form. We provide a proof for our elliptic operator in divergence form for the reader’s convenience.
@@
(C.4) $|v(0,t)|\leq Ct\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}$
This is proved by the comparison principle. We construct a barrier function in the following way. Let $M=\|v\|_{L^{\infty}(\Phi^{-1}(\Omega\cap B_{1}))}$. Let $w$ be a barrier function satisfying $L_{\Phi}(w)=0$ in $\{t>0\}\cap\Phi^{-1}(B_{1})$, $w=0$ on $\{t=0\}\cap\Phi^{-1}(B_{1})$ and $w=M$ on $\{t>0\}\cap\partial\Phi^{-1}(B_{1})$. Clearly, $0\leq w\leq M$ in $\{t>0\}\cap\Phi^{-1}(\Phi)$. The key fact we will use from the convexity of $\phi(x)$ is that $\{t>\phi(x)\}\cap\Phi^{-1}(B_{1})\subset\{t>0\}\cap\Phi^{-1}(B_{1})$ (due to the fact $\phi(x)\geq 0$). Therefore, we have $|v|=0\leq w$ on the convex part of the boundary $\{t=\phi(x)\}\cap\Phi^{-1}(B_{1})$ and $|v|\leq M=w$ on the rest of the boundary $\{t>\phi(x)\}\cap\partial\Phi^{-1}(B_{1})$. By the comparison principle, we have $|v|\leq w$ in $\{t>\phi(x)\}\cap\Phi^{-1}(B_{1})$. By the Schauder estimate of $w$ over a flat boundary $\{t=0\}$, we have $|w(x,t)|\leq CtM$ for any $(x,t)\in\{t>0\}\cap\Phi^{-1}(B_{3/4})$. This particularly implies (C.4).
Finally, by applying the previous argument for (C.4) to all the boundary points on $\{t=\phi(x)\}\cap\Phi^{-1}(B_{3/4})$, we actually have $|v(X)|\leq C\delta(X)M$, where $\delta(X)$ is the distance from $X$ to the boundary $\{t=\phi(x)\}$. Hence, by the interior Schauder estimate, for any $X\in\{t>\phi(x)\}\cap\Phi^{-1}(B_{1/2})$, we have
@@
This ends the proof of (C.3). ∎
Remark C.5.
@@
(C.7) $\left(\frac{1}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k(y)^{q}d\sigma(y)\right)^{1/q}\leq\frac{C}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k(y)d\sigma(y)=C\frac{\omega(\Delta(X,r))}{\sigma(X,r)}$
@@
by Lemma 2.11, where $G$ is the Green function for $\Delta$ with pole at infinity. Apply Lemma C.1 to $u=G$ in $B(X,r_{0})\cap\Omega$, we get that (for $s>0$ sufficiently small)
(C.8) $\frac{G(y^{\prime},s+\phi(y^{\prime}))}{s}\leq\frac{C\left\|G\right\|_{L^{\infty}(B(X,r)\cap\Omega)}}{r},$
@@
for all $r>r_{0}$. Combing the local estimate with the large-scale Lipschitz estimate, we obtain (C.8) for all $r>s>0$. By the Bourgain estimate (see e.g. [JK82a, Lemma 4.4]), one has $\left\|G\right\|_{L^{\infty}(B(X,2r)\cap\Omega)}\leq G(A_{X,r})$, where $A_{X,r}$ is a corkscrew point relative to $X$ and $r$, that is, $A_{X,r}\in B(X,Mr)\cap\Omega$ satisfies $\operatorname{dist}(A_{X,r},\partial\Omega)\approx|A_{X,r}-X|\approx r$. By Lemma 2.11 again, we have that $\frac{G(A_{X,r})}{r}\approx\frac{\omega(\Delta(X,r))}{\sigma(X,r)}$. Putting everything together, one has
@@
(C.9) $\left(\frac{1}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k^{X_{0}}(y)^{q}d\sigma(y)\right)^{1/q}\leq\frac{C}{\sigma(\Delta(X,r))}\int_{\Delta(X,r)}k^{X_{0}}(y)d\sigma(y)=C\frac{\omega^{X_{0}}(\Delta(X,r))}{\sigma(X,r)}$
@@
where $g_{X}$ is convex with Lipschitz constant at most $M$. By a covering argument, it suffices to show (C.9) for $0<r<\min\left\{r_{0}/4,\operatorname{dist}(X_{0},\partial\Omega)/4\right\}$. Since for any $X\in\partial\Omega$ and $0<r<\min\left\{r_{0}/4,\operatorname{dist}(X_{0},\partial\Omega)/4\right\}$, $G(\cdot,X_{0})$ is a solution to $\Delta u=0$ in $B(X,2r)\cap\Omega$ with $G(\cdot,X_{0})=0$ on $\partial\Omega$, one can apply (C.1) to $G(\cdot,X_{0})$ in $B(X,r)\cap\Omega$ and the rest of the argument goes as in Case 1. ∎
Remark C.10.
In the case of an unbounded Lipschitz graph domain considered in Theorem C.6, if $\phi_{1}=0$, then $\Omega$ is convex. Then $\Phi_{x_{0}}$ can be taken as identity transform and $g_{x_{0}}(x)=\phi_{0}(x)$ for any $x_{0}\in\partial\Omega$.
Theorem C.11.
Under the same assumptions as Theorem (C.6), the $L^{p}$ Regularity problem $(R)_{p}$ for the Laplace operator is solvable in $\Omega$ for any $1<p<\infty$.